OpenStudy (animecore1):

Can someone help me with this question? Solve the system. -3x + 4y = -22 2x + 7y = 5 A) (6, -1) B) (-1, 6) C) (-6, 1) D) (-1, -6)

OpenStudy (baru):

multiply first equation by 2, what do you get?

OpenStudy (animecore1):

where do I multiply the two on the first equation?

OpenStudy (baru):

on both sides of the equation. you have to multiply each term on both sides by 2

OpenStudy (anonymous):

B

OpenStudy (baru):

@FINYA direct answers are against the rules

OpenStudy (animecore1):

-6x+8y=-22?

OpenStudy (baru):

yes exactly :)

OpenStudy (baru):

in the same way, multiply second equation by 3

OpenStudy (animecore1):

could you explain why you multiply by 2?

OpenStudy (anonymous):

HMMMM sorry I'm russian and I do not know your rules)

OpenStudy (baru):

i'll show you, for now multiply 2nd equation by 3

OpenStudy (animecore1):

6x+21y=5

OpenStudy (baru):

oh sorry, i dint notice, you have forgotten to multiply the right side of the equations

OpenStudy (baru):

you have to multiply both sides of the equation

OpenStudy (animecore1):

-6x+8y=-44? and 6x+21y=15

OpenStudy (baru):

yes!! :)

OpenStudy (anonymous):

o_O Just make this 2x+7y=5 2x=5-7y x=(5-7y)/2 -3*((5-7y)/2)+4y=-22 continue and find y

OpenStudy (baru):

the rule for equations is that you are allowed to anything to them, but you have to do them for BOTH sides. if you do that, the equation will continue to hold true.

OpenStudy (animecore1):

but why are we multiplying them?

OpenStudy (baru):

so your two equations are -6x + 8y= -44 6x +21y = 15

OpenStudy (baru):

now add them , what do you get?

OpenStudy (animecore1):

29y=-29

OpenStudy (baru):

haha! excellent

OpenStudy (baru):

now can you solve and find 'y' ?

OpenStudy (baru):

hint: divide both sides by 29

OpenStudy (animecore1):

y=-1

OpenStudy (baru):

Exactly !!

OpenStudy (animecore1):

now how do I find x?

OpenStudy (baru):

awesome, so you understand x and y are unknowns, and that you need to find them

OpenStudy (anonymous):

2x+7y=5 yoy know that y = -1 so

OpenStudy (baru):

just take any 1 of you equations for example 2x + 7y = 5

OpenStudy (animecore1):

yes, I also know the substitution method but I need to know the elimination method

OpenStudy (animecore1):

then subtitute -1 for y?

OpenStudy (baru):

lol, i have very little to teach you

OpenStudy (baru):

absolutely correct

OpenStudy (baru):

what do you get for x?

OpenStudy (animecore1):

x=6

OpenStudy (anonymous):

CONGRATULATION!

OpenStudy (animecore1):

but im still confused on why u multiply by 2 and 3?

OpenStudy (anonymous):

kx=kx => 6x=6x so you can reduce it sorry for my bad english

OpenStudy (anonymous):

no?

OpenStudy (baru):

congrats indeed animecore

OpenStudy (animecore1):

could you help me with one more question?

OpenStudy (baru):

yes finya is correct. why I did that is, if both the 'x' have the same number, when you add them, they conveniently cancel and leave only y.

OpenStudy (baru):

we got 6x and -6x when we add them, we get zero :)

OpenStudy (animecore1):

oh ok

OpenStudy (baru):

i have to go now, perhaps @FINYA will help. or you can tag me, and i'll have a look when i have time

OpenStudy (animecore1):

ok thanks for your help

OpenStudy (animecore1):

could you help me @FINYA ? its a simple question but I dont know where to start on it

OpenStudy (anonymous):

I am going to english corses So I have several minutes

OpenStudy (animecore1):

ok